# Non-minimal light-curvature couplings and black-hole imaging

### v1

@@ VI Conclusions
In this work we have explored non-minimal light-curvature couplings and how they affect black-hole images, in particular the $n=1$ lensing band of the Schwarzschild black hole. Following Ref. [1], see also [79], we used that non-minimally coupled photons propagate along geodesics of effective, polarization-dependent metrics. This description holds within a finite range of values of the non-minimal coupling $\alpha$.
We are motivated by the capabilities of VLBI arrays to constrain deviations from the Kerr (or more specifically Schwarzschild) spacetime, which has led to a large body of literature [37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 32, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78]. Constraints are placed on deviation parameters in the spacetime metric or couplings in Lagrangians. These constraints bound dimensionless ratios of deviation parameters/couplings to the gravitational radius which are of $\mathcal{O}(1)$. In a standard EFT setting, one expects that the same scale suppresses all higher-order interactions. Thus, a treatment consistent with this expectation also has to account for non-minimal couplings. The specific coupling we study here is already constrained from solar-system tests and neutron-star observations. If one takes these constraints at face value [^6] In other astrophysical environments, such as, e.g., around neutron stars, other non-minimal couplings, e.g., $RF_{\mu\nu}F^{\mu\nu}$, may also play a role. The translation of observations into bounds on couplings is therefore subtle and depends on assumptions about which couplings are present in the Lagrangian; i.e., the astrophysical constraints are model-dependent in that sense., the coupling values we consider are already ruled out for supermassive black holes. However, we consider this coupling as a paradigmatic example, which creates effects in black-hole lensing bands which are likely to generalize across many other models with non-minimal couplings.
We compared the $n=1$ lensing band of a Schwarzschild black hole with and without non-minimal coupling $\alpha$. For $\alpha/M^{2}\sim\mathcal{O}(1)$, the resulting deformations of the lensing bands can become large for both photon polarizations. We find the following two results: First, in the most extreme cases within the regime of validity of the effective-metric description, the $n=1$ lensing bands with and without non-minimal coupling do not overlap. Second, lensing bands of the two polarizations differ from each other. From these results, we conclude the following:
@@
    Qualitatively, non-minimal couplings and deformations of the spacetime geometry can result in similar observational signatures, namely deformed lensing bands. There is thus a possibility that effects are degenerate and/or modifications of the spacetime geometry and non-minimal couplings may compensate to produce GR-like images, even though the underlying theory is very different from GR.
@@
    Given an external mass measurement, e.g., from stellar orbits [130] (and a spin measurement in the more realistic Kerr case), which fixes the expected photon-ring diameter in GR without non-minimal couplings, and a measurement of the photon-ring diameter, we can rule out a range of values of $\alpha$, under the assumption that the underlying spacetime geometry is that of GR.
@@
As an avenue for future research, it would be interesting to generalize our results to rotating black holes, since astrophysical supermassive black holes are expected to have non-negligible spin. The effective metrics that photons of different polarizations follow in this case have been derived in Ref. [81]. Interestingly, those effective metrics break circularity symmetry [131, 132, 133], which may occur for black-hole spacetime metrics in scenarios beyond GR [134, 135] like modified gravity and quantum gravity, e.g., [136, 137, 138, 139, 140, 141].
@@
Finally, going beyond the Horndeski coupling we considered here, there are additional non-minimal couplings not yet constrained observationally, which may have similar effects.
Ultimately, measurements of black-hole photon rings, as proposed in [85, 121], may result in constraints on various non-minimal couplings. Because these are also constrained (in combination with higher-order field-strength couplings) from positivity bounds, such constraints may provide indirect information on the properties of a more fundamental theory of gravity and matter.

### v2

@@ VI Conclusions
In this work we have explored non-minimal light-curvature couplings and how they affect black-hole images, in particular the $n=1$ lensing band of the Schwarzschild black hole. Following Ref. [Drummond:1979pp], see also [Chen:2015cpa], we used that non-minimally coupled photons propagate along geodesics of effective, polarization-dependent metrics. This description holds within a finite range of values of the non-minimal coupling $\alpha$.
[…]
We compared the $n=1$ lensing band of a Schwarzschild black hole with and without non-minimal coupling $\alpha$. For $\alpha/M^{2}\sim\mathcal{O}(1)$, the resulting deformations of lensing bands can become large for both photon polarizations. We find the following two results: First, in the most extreme cases within the regime of validity of the effective-metric description, the $n=1$ lensing bands with and without non-minimal coupling do not overlap. Second, lensing bands of the two polarizations differ from each other. From these results, we conclude the following:
@@
    Qualitatively, non-minimal couplings and deformations of the spacetime geometry can result in similar observational signatures, namely deformed lensing bands. There is thus a possibility that those effects are degenerate and/or modifications of the spacetime geometry and non-minimal couplings may compensate to produce GR-like images, even though the underlying theory is very different from GR.
@@
    Given an external mass measurement, e.g., from stellar orbits [Gebhardt:2011yw] (and a spin measurement in the more realistic Kerr case), which fixes the expected photon-ring diameter in GR without non-minimal couplings, and a measurement of the photon-ring diameter, we can rule out a range of values of $\alpha$, under the assumption that the underlying spacetime geometry is that of GR.
@@
As an avenue for future research, it would be interesting to generalize our results to rotating black holes, since astrophysical supermassive black holes are expected to have non-negligible spin. The effective metrics that photons of different polarizations follow in this case have been derived in Ref. [Chen:2023wna]. Interestingly, those effective metrics break circularity symmetry [Papapetrou:1966zz, Carter:1973rla, Carter:2009nex], which may occur for black-hole spacetime metrics in scenarios beyond GR [Delaporte:2022acp, Babichev:2025szb] like modified gravity and quantum gravity [Minamitsuji:2020jvf, Anson:2020trg, BenAchour:2020fgy, Adam:2021vsk, Eichhorn:2021iwq, Fernandes:2023vux].
@@
Finally, going beyond the Horndeski coupling we considered here, there are additional non-minimal couplings not yet constrained observationally, which may have similar effects. These couplings are, in general, not yet constrained by any observations; therefore, constraints from observations of supermassive black holes may even be of quantitative interest.
Ultimately, measurements of black-hole photon rings, as proposed in [Johnson:2024ttr, Lupsasca:2024xhq], may result in constraints on various non-minimal couplings. Because these are also constrained (in combination with higher-order field-strength couplings) from positivity bounds, such constraints may provide indirect information on the properties of a more fundamental theory of gravity and matter.